OpenStudy (anonymous):

Subtract these polynomials. (4x2 - 2x + 7) - (x2 + 3x - 4) =

OpenStudy (jlastino):

Subtract terms with like variables (4x^2-2x^2) (-2x-3x) (7-(-4)) then combine

OpenStudy (anonymous):

3x2 - 5x + 3

OpenStudy (anonymous):

you should know these too: \[\LARGE (-)\cdot (-)=(+)\] \[\LARGE (+)\cdot (+)=(+)\] \[\LARGE (+)\cdot (-)=(-)\] \[\LARGE (-)\cdot (+)=(-)\]

OpenStudy (jlastino):

Check your answer... 4-2 = ? and (7-(-4)) would be equal to 7+4 as kreshnik said